pooja195 (pooja195):

The measure of two angles in a triangle are given. Find the measure of the third angle. 30. 17° and 16.4° [A] 136.6° [B] 359.4° [C] 326.6° [D] 146.6°

pooja195 (pooja195):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

all angles in a triangle add to 180 degrees

jimthompson5910 (jim_thompson5910):

so if x is the missing angle, and the other two angles are 17 and 16.4, then 17+16.4 + x = 180

jimthompson5910 (jim_thompson5910):

and you would solve for x

pooja195 (pooja195):

would i add?

jimthompson5910 (jim_thompson5910):

at first, yes 17+16.4 + x = 180 33.4 + x = 180 ... add

jimthompson5910 (jim_thompson5910):

then you would subtract 33.4 from both sides to isolate x

pooja195 (pooja195):

so my answer woood be 146.6 right

jimthompson5910 (jim_thompson5910):

yep

pooja195 (pooja195):

thank you!!

jimthompson5910 (jim_thompson5910):

np